Solve the following quadratic functions
(a) Y=x2-7x+12
(b) Y=3x2-7x-6

Answer by stanbon(75887)

Solve the following quadratic functions
(a) Y=x2-7x+12
y = (x-4)(x-3)
====---------=====
(b) Y=3x2-7x-6
y = (3x^2-9x+2x-6)
y = 3x(x-3)+2(x-3)
y = (x-3)(3x+2)
=====================
Cheers,
Stan H.
=================